OpenStudy (meilendurcer):

Which set of line segments with the given lengths form a right triangle? A. 5, 12, 17 B. 7, 11, 15 C. 8, 15, 17 D. 14, 16, 20

HanAkoSolo (jamierox4ev3r):

you would test all of these side lengths by plugging it in to the pythagorean theorem

HanAkoSolo (jamierox4ev3r):

\[A^2+B^2=C^2\]

HanAkoSolo (jamierox4ev3r):

a and b are the side lengths and c is the hypotenuse

HanAkoSolo (jamierox4ev3r):

is this making sense so far @meilendurcer ?

OpenStudy (meilendurcer):

@Jamierox4ev3r yes

HanAkoSolo (jamierox4ev3r):

awesome, so you plug in your numbers \[5^2+12^2=17^2\]

HanAkoSolo (jamierox4ev3r):

\[25+144\neq289\]

HanAkoSolo (jamierox4ev3r):

so we know that A is not correct. Can you prove the rest based from what i did with the first?

HanAkoSolo (jamierox4ev3r):

@meilendurcer ?

OpenStudy (meilendurcer):

I can try

OpenStudy (meilendurcer):

\[7^{2}+11^{2}=15^{2}\]

HanAkoSolo (jamierox4ev3r):

srry i was afk @meilendurcer \[7^2+11^2=15^2 =49+121\neq225\] so we know that B. is not correct either, which leaves us with either C. or D.